### v2

# Detection and Tracking of MAVs Using a Rosette Scanning Pattern LiDAR

Abstract
The use of commercial Micro Aerial Vehicles (MAVs) has surged in the past decade, offering societal benefits but also raising risks such as airspace violations and privacy concerns. Due to the increased security risks, the development of autonomous drone detection and tracking systems has become a priority. In this study, we tackle this challenge, by using non-repetitive rosette scanning pattern LiDARs, particularly focusing on increasing the detection distance by leveraging the characteristics of the sensor. The presented method utilizes a particle filter with a velocity component for the detection and tracking of the drone, which offers added re-detection capability. A Pan-Tilt platform is utilized to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense. The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments. Our approach achieved accuracy on par with the state-of-the-art indoor method while increasing the maximum detection range by approximately 80% beyond the state-of-the-art outdoor method.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I INTRODUCTION
Micro Air Vehicles (MAVs) have contributed to the development of many fields, such as precision agriculture [1], modern production [2], transport industry [3] and mining industry [4]. In addition to the above, the utilization of drones (in this paper the phrases drone and MAV are interchangeable) for hobby purposes is also developing explosively, because they have become affordable and easy to handle [5], increasing the risk of inappropriate use. Hobby drones can be used for unauthorized recording, violating people’s right to undisturbed privacy [6, 7]. Furthermore, they can also threaten protected buildings (e.g. airports, factories, power plants) [8]. Due to the increased security risks, the development of autonomous drone detection and tracking systems has become a priority area of research.
In this study, we tackle the challenge of drone detection and tracking using non-repetitive rosette scanning pattern LiDARs, particularly focusing on increasing the detection distance by leveraging the characteristics of the sensor.
We use the rosette scanning pattern Livox AVIA LiDAR sensor, that uses multiple, aligned laser beams. The beams move in a non-repetitive rosette-like pattern inside the sensor, that covers the entire Field of View (FoV) by using Risley prisms [9]. This type of sensor design results in faster measurements, and cheaper devices with fewer moving components compared to the more common rotating models. An important feature of this type of LiDAR is that it captures more measurement data from a target object in the center of the sensor’s FoV, as the laser beams pass through the center each time they travel along a rotating 8-shaped pattern.
Several sensors can be used for drone detection, such as radar [10, 11], LiDAR [12, 13], camera [14, 15, 16], radio-frequency [17, 18] and acoustic sensors [19, 20]. The detection efficiency can be further improved by using sensor fusion [21, 22]. However, each sensor modality should also give good results when used independently [23].
The state-of-the-art drone tracking algorithms using LiDARs either use a classical rotating LiDAR [13] or place a rosette scanning pattern sensor on a stationary mount [24]. Dogru et al.[13] achieve a detection distance of $70\;m$ outside using a Velodyne VLP-16444https://velodynelidar.com/products/puck/. This detection distance can only be extended to $100\;m$ by adding a retroreflector to the drone. Catalano et al.[24] use a LiVOX Horizon [^5] https://www.livoxtech.com/ to achieve a few $cm$ accuracy using a method based on Kalman Filter which has no re-detection capability in case of lost tracking. Their method was only tested indoors so there is no information on the maximum detection distance.
Our proposed algorithm uses a particle filter with a velocity component to achieve fast detection and tracking with added re-detection capability. Leveraging the point distribution of the LiVOX AVIA LiDAR [^6] https://www.livoxtech.com/avia, the proposed method keeps the MAV in the middle of the sensor’s FoV, where the measurements are most dense, using a pan-tilt platform. With this approach we achieved accuracy on par with the state-of-the-art indoor method using a similar sensor [24], while increasing the maximum detection range approximately $80\%$ beyond the state-of-the-art outdoor method [13] by detecting drones up to $120m$ away without attached retroreflector. In addition, the number of points reflected from the drone at the same distance increased several folds over the full range.
To focus on this challenge, we made two assumptions: (i) the background of the detection area is assumed to be static, and (ii) we disregard the classification problem and assume that all moving objects in the area are drones.
To summarize, this paper enhances the state-of-the-art in LiDAR-based drone tracking with the following contributions:
  • A novel MAV tracking approach for LiDAR utilizing a particle filter with a velocity component is proposed. The updating of particle states of the filter is straightforward using the LiDAR point clouds and the proposed solution offers added re-detection capabilities.
  • Our method uses a Pan-Tilt platform to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense.
  • The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments [^7] Video demonstration of experiments at https://youtu.be/VdePm3on8vo.
The remainder of the paper is organized as follows: related work is presented in section II, the proposed algorithm in section III, the experiments in sections IV and V, and finally the conclusion in section VI.
II RELATED WORK
Tracking MAVs is challenging due to their small size and rapid movement. Traditionally, vision-based approaches have dominated this field. Methods leveraging neural networks, specifically YOLO [25], are widely used for detecting and classifying objects in images and videos. These methods are highly effective in controlled environments with good visibility, as shown in works like [14, 15, 16, 26]. Vision-based techniques leverage large datasets and complex neural network architectures to improve accuracy, but they often struggle in outdoor scenarios where lighting and occlusions affect visibility. Furthermore, the maximum detection distance is not only limited by the resolution of the camera but also the input resolution of the neural network used. In a recent work [27], the maximum detection range of such a system is 40 m.
In response to these limitations, there is a growing interest in non-camera-based solutions, particularly LiDAR, which provides accurate 3D spatial information. LiDAR offers a distinct advantage in detecting and tracking MAVs in low-visibility conditions, where cameras are less effective.
Model-based object detection approaches for LiDAR, such as [28], PointNet [29], PointNet++ [30], VoxelNet [31], and PointPillars [32] are used to detect larger objects like cars or humans, thus making them insufficient for the detection and tracking of drones.
Unlike model-based object detection, model-free methods detect dynamic objects by comparing consecutive scans, making them more versatile. For instance, [33] tracks moving objects by identifying motion patterns between scans, while [34] employs a Kalman filter combined with Iterative Closest Point (ICP) alignment to track objects in dense 3D LiDAR data. Additionally, [35] introduces a method that clusters point clouds using a region growth algorithm and filters objects based on geometric features such as height range and maximum diagonal width. These methods utilize classical rotating LiDARs and also track larger objects like cars or humans.
In [36] a method is proposed for finding MAVs in a point cloud generated by a rosette scanning LiDAR on a moving robot. This approach is further improved by adding an Kalman Filter based tracking method and extensive evaluation in [24]. In these papers, the integration time of the LiDAR is dynamically adjusted to adapt to the speed and distance of the target’s movement enabling the tracking of MAVs indoors up to about $30\;m$. However, this method does not allow for regaining tracking once lost. Conversely, our aim is to align the rosette scanning pattern LiDAR to ensure continuous centering of the target drone within the sensor, thereby maximizing the number of returned points. Additionally, our approach incorporates a robust tracking method capable of regaining tracking.
[…]
LiDAR sensor performance in outdoor environments is highly influenced by weather conditions [37], [38]. According to [39] the detection range of LiDAR sensors is limited to $25\;m$ in foggy conditions when the meteorological visibility is under $40\;m$. Based on this a significant reduction in the detection distance and the number of detected points can be expected for outdoor tests under foggy weather conditions. While our experiments confirmed this trend, our method could still track the drone to a little more than $50\;m$.
III THE PROPOSED ALGORITHM
The flowchart of our method is presented in Fig. 1, where the main steps of the algorithm are shown. Note on the figure, when the system is installed at a new location, the first step is to build a background model (see Section III-A). This is used to filter the static background from the point cloud to detect the target object.
Figure 1: Overview of the proposed method.
Next, as part of the object tracking, the particle filter[40] performs the prediction step. If there is a new incoming measurement point cloud, the background model is subtracted to reduce static noise (details are given in Section III-B). The three steps of the particle filter (predict, update, and resample) are explained in detail in Section III-C.
Once detected, the drone is kept in the center of the LiDAR’s FoV by orienting the pan-tilt turret towards the position computed by the particle filter (see Section III-D). As a result, the measured point density reflected from the drone is maximized.
III-A Building the Background Model
In our application, the background is considered static throughout the drone tracking, hence it is built once at the beginning of the algorithm. Consequently, the proposed method considers everything that is not recorded in the background a drone. The turret is set to scanning mode and performs movements to scan a pre-defined area. Because of the LiDAR’s non-repetitive way of scanning, the point density of the whole scene will increase the longer the scanning time is.
The background model is represented with OctoMap[41], which is an efficient probabilistic 3D mapping framework using octrees. The octree tree structure results in fast access times when querying individual voxels.
Based on the position and orientation of the LiDAR, the incoming point cloud is transformed into the static coordinate system of the background model. The points are filtered based on position thresholds; the ground plane, near points, and far away points are removed. After filtering, the point cloud is inserted into the OctoMap of the background model. As the last step of building the background model, after inserting all the collected point clouds, inflation is performed to account for measurement noise (e.g. vibrations of the turret base) as well as small changes in background objects (such as leaves moving on the trees).
III-B Processing New Measurements
The point measurements are received in the Cartesian frame of the LiDAR and transformed into the World frame based on the turret’s pan and tilt angles. Then incoming point cloud is filtered as during the background building: the ground plane, near points, and far away points are removed. And finally, the background model is subtracted. There are two additional filters to remove the reflected points that are not close enough to each other and cannot belong to the target object. The filters are the RadiusOutlierRemoval and StatisticalOutlierRemoval algorithms from the Point Cloud Library (PCL) [42].
III-C Object Tracking Using Particle Filter
The drone’s position is estimated using a particle filter, which is a recursive Bayesian state estimator that uses discrete particles to approximate the posterior distribution of the estimated state. It accommodates both linear and non-linear process and measurement models, allowing for flexible, arbitrarily shaped probability distributions. This capability is especially useful for MAVs, which exhibit a wide range of motion patterns with abrupt changes in direction and speed across all three dimensions. Additionally, LiDAR measurements enable straightforward updates of the particle states.
The particle filter operates in three main steps: prediction, update, and resampling. The prediction step uses the current state and transition model to predict the next state. In the update step, incoming filtered point cloud data refines the state estimate via Bayes’ theorem. Finally, resampling aligns the particles with the posterior distribution, addressing the issue of particle degeneracy.
The particle filter is implemented following Algorithm 2. in [43]. Each particle represents a possible state of the drone, where the state is the 3D location and 3D velocity of the drone ($s=[x,y,z,v_{x},v_{y},v_{z}]^{T}$).
III-C1 Predict
First, the position of each particle is predicted based on its previous position, velocity, and added process noise. The prediction noise $p_{\{x,y,z\},k}$ is sampled from a 00 mean Gaussian distribution with $\sigma_{pred}$ standard deviation and $T$ denotes the period of the particle filter.
$p_{\{x,y,z\},k}\sim\mathcal{N}(0,\sigma_{pred}^{2})$ (1)
$\begin{bmatrix}x_{k}\\ y_{k}\\ z_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}+v_{x,k-1}\cdot T+p_{x,k}\\ y_{k-1}+v_{y,k-1}\cdot T+p_{y,k}\\ z_{k-1}+v_{z,k-1}\cdot T+p_{z,k}\end{bmatrix}$ (2)
The velocities are computed only if the drone is tracked. The tracking can be considered stable when the standard deviation of the particle positions ($\sigma_{particles}$) is around the parameter $\sigma_{pred}$. The $\sigma_{particles}$ value increases with $\sigma_{pred}$ plus the additional effect of the velocity in every iteration when the drone is lost. The drone is considered tracked if it is under $\sigma_{threshold}=2\cdot\sigma_{pred}$. In this case, the velocities are predicted, or else they are set to 00.
$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}(x_{k}-x_{k-1})/T\\ (y_{k}-y_{k-1})/T\\ (z_{k}-z_{k-1})/T\end{bmatrix}$ (3)
III-C2 Update
The weight of each particle is updated by the product of the previous weight and the Gaussian distribution with $\sigma_{m}$ standard deviation and 00 mean evaluated at $d_{i}$, where $\sigma_{m}$ represents the measurement noise and $d_{i}$ is the distance from the $i$th particle to the closest LiDAR measurement point. This product is normalized, so the sum of $w_{i}$ will add up to $1$, where $n$ is the number of particles.
$w_{i,k}=\frac{w_{i,k-1}\cdot\mathcal{N}(d_{i};0,\sigma_{m}^{2})}{\sum_{i}^{n}w_{i}}$ (4)
III-C3 Resample
After every update step, the particles are resampled using their weights as the probability of being drawn for the next iteration.
III-C4 Particle initialization and reset
The particles are initialized randomly in a cylinder, whose parameters (3D location, radius, and height) are set based on the detection area. The particles can be reset to a new cylinder during runtime, thus giving the operator the capability to reset the tracking to another object. During tracking $\sigma_{particles}<\sigma_{threshold}$ which means 2 objects are only mistaken if they are closer than $\sigma_{threshold}$ which is almost certainly a collision of the objects.
III-D Turret control
Turret control is done in two ways: in initialization or tracking mode. During initialization mode, the pan-tilt turret collects data points for a specified time in a predefined area to build the background model. In tracking mode, the control commands are calculated based on the current position of the turret and the estimated position of the drone. The tracking mode aims to keep the drone in the middle of the FoV to maximize the density of reflected points from the drone. In case there are multiple drones in the FoV, the algorithm follows the one closest to the location of the particles, which can be reset by the operator.
III-E Implementation details
The Robot Operating System (ROS) was used to integrate and manage communication between program components, offering a reliable framework for modular nodes. Three nodes were implemented: a LiDAR wrapper node to fix point cloud conversion issues, a pan-tilt turret wrapper node to translate MAV position estimates into turret commands, and a MAV detection and tracking node that holds the core algorithm, including the particle filter.
IV INDOOR EXPERIMENTS and EVALUATION
First, the experimental results are presented and examined, starting with the initial detection of the MAV following the construction of the background model. Next, the core of the algorithm, the tracking process, is evaluated in terms of accuracy. Finally, the re-detection capability after tracking loss is assessed.
The experiments were conducted within the MIMO arena [26] within the HUN-REN SZTAKI Institute for Computer Science and Control incorporating an OptiTrack motion capture system in a $7\times 8\;m^{2}$ office space with a height of $3.5\;m$. Further hardware includes MAVs (Crazyflie 2.0 by Bitcraze), the turret (WidowX Dual XM430 Pan & Tilt turret by Trossen Robotics), and the rosette scanning pattern LiDAR (Avia LiDAR sensor by Livox) with an integration time of $100\;ms$ ($10\;Hz$).
In all our experiments the particle filter was run at $15\;Hz$ frequency which assured the processing of all LiDAR measurements. The processing ran on an 8th-generation desktop Intel i7 processor with less than 30% utilization.
IV-A Detecting the MAV
Figure 2: Initial detection error and $\sigma_{particles}$ of the particle filter. The particle filter locks onto the drone almost instantaneously after receiving the first measurement at around 2 seconds.
Fig. 2 shows the error as the distance of the estimated state to the true state in meters of a representative detection. The average of the standard deviation of the particles in each dimension is also plotted (denoted as $\sigma_{particles}$).
When there is no incoming measurement only the predict step of the particle filter is performed. In all of our experiments $\sigma_{pred}=0.1\;m$ and $n=500$, which turned out to be a good compromise between being able to track fast-moving drones and providing accurate measurements.
While the drone is out of the FoV of the LiDAR sensor, the particles remain idle, only influenced by the added noise in the predict step which results in a slowly inflating particle cloud with a slowly rising $\sigma_{particles}$. In this case, the MAV is resting on the floor and, because of its small size, is hidden inside the background model. As the drone starts hovering and measurements of it arrive at around 2 s the particles gather and allow for the first good estimation. Parallel with this the $\sigma_{particles}$ of the estimation drops sharply. This means the drone is tracked and the particle filter also predicts velocity.
IV-B Tracking the MAV
Figure 3: Different drone testing flight patterns with different colors (vertical, horizontal, fast, and ’lost and found’). The outlined circles show the starting position of each trajectory while the arrows denote the moving direction of the drone. The dashed part of the arrow means that the MAV view is obstructed from the LiDAR’s point of view by an obstacle. The size of each trajectory is in the legend of the figure.
For calculating the accuracy, the MAV was flown following certain flight patterns (described in Fig. 3) in a repeated manner. The ground truth data was obtained by the motion capture system within the MIMO arena by recording the pose of the LiVOX sensor and the position of the drone and aligning the LiVOX measurements with the MIMO arena using the recorded pose. The patterns in Fig. 3 were drawn from the point of view of the LiDAR sensor where the middle of each trajectory in the horizontal direction was in front of the sensor with a minimum distance of around $3\;m$. Each circle represents a stationary wait time of $2\;s$ and the one with a black outline represents the first position of the respective pattern. The trajectory represented by the arrows in-between was flown in $3\;s$ in case of the vertical, horizontal, and ’lost and found’ pattern which resulted in a maximum speed of about $0.6\;m/s$, while it was flown in $2.25\;s$ in case of the fast pattern which resulted in a maximum speed of around $1.2\;m/s$. In each experiment, the drone flies to the first point and then it flies the pattern 3 times in the case of the vertical and horizontal patterns and 4 times in the case of the others.
In Fig. 4, the error of the particle filter is plotted with the ground truth speed calculated from the MIMO measurements. It is visible that the tracking is stable over time with an average position error of under $5\;cm$ in the case of the horizontal and vertical measurements while it is around $7\;cm$ for the fast pattern.
Figure 4: Position error of the particle filter plotted together with ground truth speed from the MIMO arena for the vertical (top), horizontal (middle), and fast (bottom) patterns. The mean ($\mu_{error}$) and standard deviation (±$\pm\sigma_{error}$) of the particle filter error are also plotted.
It can be seen on all the plots of Fig. 4, that the error of the particle filter is dependent on the speed. When the drone is at a waypoint, not moving, the error of the filter is around $3$ to $5\;cm$, while at the peak speeds of $1.2\;m/s$ of the fast pattern (not counting the one outlier), the error is around $20\;cm$. This dynamic difference of about $15\;cm$ in the error can mostly be explained by the different delays in the pipeline. The internal processing time of the LiDAR firmware is $2\;ms$ based on the data-sheet, the LiDAR scan integration time is $100\;ms$, and there is some additional overhead in the formatting, stamping, and publishing of the data to ROS topics. The background subtraction and the application of filters took $10.9\;ms$ with a $\sigma=16.8\;ms$ standard deviation on the fast trajectory and the predict, update and resample steps took a combined red1.$1.11\;ms$ with a $\sigma=1.12\;ms$. The whole delay is about $120\;ms$, in which the drone travels roughly $14.4\;cm$. From the elapsed time, the computation takes only around $20\;ms$ resulting in convenient real-time operation.
The particle filter’s inherent functionality causes the particles to consistently lag slightly behind the measurements. Additionally, the relative motion between the drone and the LiDAR sensor naturally introduces smearing in the point cloud, accounting for the remaining dynamic error.
Overall, when the drone is stationary, the error remains minimal and consistent. However, during movement, a dynamic error arises that increases linearly with speed. This affects the tracking accuracy but does not compromise its robustness.
For comparison, in [24] the RMSE is reported as the metric of accuracy on different trajectories. From their paper, the trajectory T1 is most similar to our horizontal pattern and the trajectory T3 is most similar to our vertical case. We consider them similar because the trajectories are mostly in the horizontal and vertical planes respectively and mostly contain straight lines and turns close to right angles. Their reported best RMSE for T1 and T3 are $0.0773\;m$ and $0.0421\;m$ while the RMSE for our horizontal and vertical patterns are $0.0570\;m$ and $0.0531\;m$ respectively. While the two methods work very similarly in terms of accuracy, our particle filter approach has the added benefit of being able to regain the tracking after it was lost.
IV-C Losing and regaining Tracking
An advantage of utilizing a particle filter is that the detection and tracking properties are part of the same routine, which leads to good performance in terms of regaining tracking after the target is lost.
(a) Stable before loss
(b) Right after loss
(c) Lost for some time
(d) Stable after regain
Figure 5: Different stages of the particle cloud during tracking loss and regain. The particle cloud can be seen with yellow in front of the OctoMap of the background model. The number of particles was 500 in this experiment.
Figure 6: Particle filter error and $\sigma_{particles}$ are plotted together for a typical loss and regain event. The tracking was lost at around 4 seconds and was regained at 8 seconds.
Fig. 5 shows a typical loss and regain event of the MAV tracking from (a) to (d) while Fig. 6 shows the corresponding error and $\sigma_{particles}$ characteristics. This is one traversal of the drone through the 3 waypoints of the ’lost and found’ pattern. First, the normal MAV tracking is depicted in Fig 5(a). The particles are grouped around the drone with $\sigma_{particles}<0.2\;m$. When the drone starts moving, the error increases slightly but the tracking is still stable. At around 4.5 seconds the drone gets behind the obstacle and $\sigma_{particles}$ increases over $0.2\;m$ meaning the tracking is lost and the particle cloud starts inflating (Fig. 5(b)). Note on the plot, that at first the $\sigma_{particles}$ jumps sharply when the velocity is still estimated and then settles to a much gentler slope when only the prediction noise is added to the particle positions. The particle cloud inflates further to almost $\sigma_{particles}=0.8\;m$ which is depicted in Fig. 5(c) while the error stays constant. This is because the drone hovers behind the obstacle and the estimated pose is the mean of the particle positions which is around the position where the tracking was lost. Then the drone flies out from behind the cover and the particle update and resample steps of the filter condense the cloud instantly which is well depicted on the sharp drop of $\sigma_{particles}$ below $0.2\;m$ in Fig. 6. The tracking is stable again in Fig. 5(d) and the error also decreases to the static error at about 9 seconds when the drone stops moving.
Overall the regaining of tracking after loss is practically instantaneous allowing robust tracking even in scenarios where the trajectory is partially occluded.
V OUTDOOR FIELD TESTS and COMPARISON
After the successful indoor tests, we validated the developed method in an outdoor environment as well. We performed the measurements with the presented method under two different weather conditions (on a sunny and a foggy day). The purpose of the validation is to test the track movement and to determine the detection distances and number of detected points in outdoor conditions, on a commercial drone. Finally, we compared the performance of our method to the previous state-of-the-art [13].
V-A Outdoor test
For outdoor tests, the DJI Phantom drone was used on approximately flat terrain. DJI Phantom 4 Pro V2.0 measures $251\;mm\times 398.78\;mm\times 172.72\;mm$ and weighs $1375\;g$ with a battery and propellers attached. An important parameter for detection is the frame size (the largest distance between the motors), which is $350\;mm$ for this type of drone. The measurements were carried out in the factory state of the drone, we did not improve the detection range by using special paint or other methods. For a video of the outdoor experiment, kindly check the video attachment here: https://youtu.be/VdePm3on8vo.
Figure 7: Weather on the test days. Sunny experiment with clouds on the left and the foggy experiment with severely limited viewing conditions on the right.
A sunny day was chosen for the first and a foggy day for the second measurement day when visibility was also severely limited (see Fig.7).
V-B Results
During both tests, the drone was moving under such conditions that both horizontal and vertical distances could be analyzed. In both cases, the controlled pan-tilt turret followed the drone correctly, thus the drone tracking was successful.
(a) Sunny
(b) Foggy
Figure 8: Trajectory and speed of the drone on a sunny and foggy day. The turret was at the origin of the coordinate system.
When evaluating drone detection, an important value is the maximum detection distance. The drone’s trajectory was successfully tracked up to a distance of $128\;m$ in the sunny test environment (Fig. 8(a)), and up to a distance of $58\;m$ in the foggy test environment (Fig. 8(b)). As explained earlier in subsection IV-A, we consider the drone tracked when $\sigma_{particles}<0.2\;m$. We consider the tracking lost after 10 subsequent missing measurements, but it is important to note that we had some occasional detections at distances of up to $145\;m$ on the sunny day.
In both scenarios, we calculated maximum speeds of a little more than $6\;m/s$ from the tracking positions and timestamps.
V-C Comparison
We compare our outdoor test results with another study in which the drone detection efficiency of the VLP-16 LiDAR was tested. In the work of Dogru et al.[13], the size of the drones is larger than the one used for our measurement, which affects the number of reflected points. Furthermore in their paper, there are measurements in which a retroreflector was used to increase the number of points reflected to the sensor.
TABLE I: Comparison of maximum detection distances
Measurement, method Weather With retroreflector Distance($m$)
Our test I. Sunny No 128
Our test II. Foggy No 58
Dogru et al.[13] Sunny Yes 100
Dogru et al.[13] Sunny No 70
Our maximum distance of detection is about $1.5$ to $2$ times greater than their result as in TABLE I.
Another important aspect is the number of detected points, which greatly impacts the robustness of tracking. In sunny weather, except for the last part of the detection range, the number of detected points in the result is $3$ to $5$ times higher than in the current state-of-the-art (see Fig. 9).
Figure 9: Number of detected points depending on distance. The proposed algorithm returns considerably more points in every distance range than [13].
In the case of foggy weather, the detection distance was significantly reduced, as expected in advance based on the literature[39].
VI CONCLUSION
This article presents a novel tracking method utilizing a rosette scanning pattern LiDAR mounted on a Pan-Tilt platform. The proposed algorithm has been rigorously tested in both indoor and outdoor environments, demonstrating accuracy on par with the current state-of-the-art indoor method [24] and considerably surpassing the leading outdoor method [13] in terms of detection range and point density. Notably, our method now tracks all dynamic objects, and ongoing work aims to further enhance the system by differentiating between drones and other flying objects.